# Deterministic generation of single B centers in hBN by one-to-one conversion from UV centers

### v1

@@ Abstract
Among the variety of quantum emitters in hexagonal boron nitride (hBN), blue-emitting color centers, or B centers, have gathered a particular interest owing to their excellent quantum optical properties. Moreover, the fact that they can be locally activated by an electron beam makes them suitable for top-down integration in photonic devices. However, in the absence of a real-time monitoring technique sensitive to individual emitters, the activation process is stochastic in the number of emitters, and its mechanism is under debate. Here, we implement an in-situ cathodoluminescence monitoring setup capable of detecting individual quantum emitters in the blue and ultraviolet (UV) range. We demonstrate that the activation of individual B centers is spatially and temporally correlated with the deactivation of individual UV centers emitting at 4.1 eV, which are ubiquitous in hBN. We then make use of the ability to detect individual B center activation events to demonstrate the controlled creation of an array with only one emitter per irradiation site. Additionally, we demonstrate a symmetric technique for heralded selective deactivation of individual emitters. Our results provide insights into the microscopic structure and activation mechanism of B centers, as well as versatile techniques for their deterministic integration.
I Introduction
@@
Optically active point defects in crystals, commonly referred to as color centers, are of great significance in quantum information science due to their ability to act as single-photon emitters (SPEs) and their compatibility with nanostructures and devices [1]. A well-known example is the nitrogen-vacancy (NV) center in diamond, which forms the foundation of many emerging quantum network architectures [2]. More recently, the discovery of color centers in van der Waals materials has opened new avenues for integration and applications in quantum technologies [3], with hexagonal boron nitride (hBN) being a particularly promising host. This wide-bandgap semiconductor can indeed integrate high quality SPEs [4, 5]. A specific class of hBN emitters is distinguished by its remarkable quantum optical properties. The so-called B centers reproducibly emit at 436 nm and exhibit photon indistinguishability at low temperature [6, 7]. They are not found natively but can be locally activated by a focused electron beam [8, 9, 10], which is an asset for the controlled integration in photonic devices [11, 12]. However, the activation mechanism is still debated and yields an unpredictable number of SPEs, due to the lack of a technique for heralding the successful creation of single emitters.
@@
The microscopic structure of the B centers is still under debate. Some important clues have been however identified. In particular, the B center activation is efficient only in crystals exhibiting the optical signature of UV centers [10]. These UV color centers emitting at 4.1 eV are commonly found in carbon-rich hBN [13, 14, 15, 16, 17, 18] and are attributed to in-plane carbon dimers (CBCN) [19, 20, 21]. Cathodoluminescence (CL) has been shown to allow for real-time monitoring of the emitter population dynamics under electron irradiation, allowing to study the kinetics of B center creation [22] and to observe the concurrence of a decrease in the population of UV centers [23], albeit with a different timescale. These studies have been performed on ensembles and at room temperature, therefore the possible relation between UV and B centers remains unclear. Moreover, the activation and deactivation of presumably individual B centers has been observed [8, 24], but without insights about the mechanism of such dynamics and the relation to individual UV centers. Concerning the microscopic structure of B centers, several candidates have been proposed, such as a split nitrogen interstitial [25, 26], a carbon tetramer [27, 28], and a vertical carbon dimer [25]. The latter hypothesis seems to have been confirmed very recently by a bundle of consistent clues obtained with combined STEM/HAADF and CL measurements [24].
@@
Here, we investigate the activation and deactivation of individual B and UV centers upon irradiation by an electron beam, through a specifically designed CL detection setup based on avalanche photodiodes (APDs), which makes the in-situ monitoring sensitive to the single emitter level in both wavelength ranges. We show that individual creation events of B centers are accompanied by a simultaneous abrupt deactivation of a UV center, correlated both in time and space. This suggests a one-to-one transformation of UV to B centers. The conversion is partially reversible, as the opposite process is also observed. At long times, B centers stabilize into a steady state population, as demonstrated in prior work [22, 23]. In a second step, we make use of the in-situ monitoring signal as a marker that heralds successful creation of individual emitters. We demonstrate a small array of emitters with strongly sub-Poissonian statistics of emitter numbers, as confirmed with the measurement of their second-order correlation function. We then show that accidental double creations can be corrected for using heralded photobleaching, further reducing the variations in emitter numbers. In a conclusive discussion, we tentatively provide a mechanism for supporting the observed UV-blue conversion, consistent with the vertical carbon dimer hypothesis for the microscopic structure of B centers.
II Interconversion of UV and B centers
@@
The experimental setup which allows for in-situ monitoring of the activation and deactivation of individual color centers in the UV and blue spectral range is presented in Figure 1a. The sample is placed inside a scanning electron microscope (SEM), where a holed parabolic mirror is positioned above the sample to collect the emitted photons outside the microscope. A spectrometer is placed at the output port of the SEM. Figure 1b shows the CL spectrum of a hBN crystal under electron irradiation. Three recognizable features can be observed. The exciton emission is visible at about 215 nm; the three equally spaced peaks in the 300-350 nm range are the signature of the UV emitter CBCN with a zero-phonon line at 305 nm and its phonon replica; the peaks at 440 nm and 465 nm originate respectively from the B center zero-phonon line and optical phonon replica [13, 29, 21]. Using a flip-mount mirror, we can channel the emission to a single-photon detection system. The emitted light is then separated into two frequency bands, UV and blue regions respectively, using a dichroic mirror and bandpass filters. The bandwidth of both detection channels is depicted by the shaded areas on Fig. 1b. Both signals are simultaneously detected by APDs. This setup allows to be sensitive to the CL signal associated with individual UV and B centers.
@@
We first select a hBN crystal of thickness 145 nm, where the density of UV centers is low enough to spatially resolve individual emitters (see Methods for sample preparation). Figure 2a shows a series of four consecutive CL scans of the same region of space. The scanning time is about ten minutes, and the scan direction is from left to right, then from top to bottom. Both UV and blue channels are recorded in parallel, simultaneously. The top (bottom) panels plots the UV (blue) band signal. We refer to these successive maps as scan 1 to 4. In both spectral ranges, various point emitters can be observed. In addition, some emitters of both species either appear or disappear during scans. To ease the discussion, we select five representative sites, depicted by circles in both spectral regions and labeled by letters A to E. The circle colors indicate that the emitters are active (green circles), inactive (red circles) or switching (two-colored circles). In the UV maps, sites A and B indicate UV emitters that are active during scans 1 and 2 (upper panel), while no blue emitter is observed at these same sites (lower panel). In scan 3, the UV emitters located sites A and B are deactivated, which shows up as interrupted spots with CL signal only in the upper part. Simultaneously, signal appears in the blue range at the same locations. The switching appears instantaneous at the timescale of the integration time per pixel (100 ms), as the blue signal appears at the same pixels where the UV signal disappears. In scan 4, no UV emitter is visible in sites A and B, while blue emitters now appear as full spots. Sites C and D exhibit the inverse process along the four scans, with blue emitters initially present in scan 1, switching to UV during scan 2, and staying visible as UV emitters in the two following scans. Finally, site E undergoes a bidirectional process, with an initial switching from UV to blue during scan 1, staying in the blue state during scans 2 and 3, and switching back to UV during scan 4. Other similar events can be observed on these CL scans, as well as all other scans we performed. We illustrate this further by showing consecutive scans of two emitters on Figure 2b and c. Both emitters exhibit multiple switching events between the UV and the blue state while the electron beam is passing over them. Additional examples of consecutive scans are shown in the Supporting Information section S1. Overall, our observations can be summarized as follows.
@@
(4) The switching occurs exactly under the electron beam, as testified by truncated CL spots, with the signal appearing or disappearing at the emitter position. This can be observed as incomplete lines of bright pixels switching while the electron beam is scanning over the SPE. This implies that the process is driven locally by the electron beam, and not by indirect consequences of the irradiation process, such as charge accumulation or heat. This also demonstrates that, in samples with high concentration of UV centers, the resolution of the B center creation is solely given by the SEM resolution.
(5) Essentially all occurrences where a single UV center disappears while scanning are associated with the appearance of a B center. This suggests that the conversion to B center is the main – if not the only – decay process of the UV center under electron irradiation. Other mechanisms were proposed [30] but are likely slower or less probable, at least in our range of irradiation parameters.
@@
III Heralded activation of B centers
We now select a flake with a higher density of UV centers, such that conversion can occur at any location. Such flakes exhibit a large UV signal, similar as the spectrum shown Fig. 1b. We continuously irradiate the crystal at a fixed position, while recording the signal of B centers with the APD. Figure 3 shows a representative example of the intensity trace measured during 80 s. Discrete increases of the signal can be observed, showing that individual creation events can be unambiguously identified. The signal associated with the creation of individual emitters is about to counts per second. A few downward steps are also visible, associated with deactivations as observed in the previous section. However, the overall trend is a global increase of the number of emitters, consistently with prior work on large ensembles [22, 23]. This suggests that the blue emitters are stabilized under continuous irradiation. Each upward step constitutes a signal heralding the success of individual B center activation.
@@ III Heralded activation of B centers
The sample is subsequently characterized in photoluminescence (PL) at room temperature, in a confocal microscope. A pulsed laser emitting at 405 nm is used to excite the emitters. Figure 4b shows a PL confocal map of the sample. Eight luminescent spots are visible. The slight misalignment of the spot array originates from hysteresis of the piezoelectric positioner used to define the irradiation sites. The emitters are characterized in a Hanbury Brown and Twiss setup to estimate the number of emitters. A value of is attributed to a single emitter. An example of is shown on Figure 4c, corresponding to the circled spot on Figure 4b. Among the eight sites, seven are single emitters, and one, indicated by the dotted circle on Figure 4b, has a higher . The latter corresponds to an event where two steps were observed before the electron beam was blanked. A workaround for such cases is proposed in the next section.
This proof of principle demonstrates the ability to control the number of activated color centers. Without the active feedback developed in this work using in-situ CL monitoring, a Poisson distribution of the emitter numbers would have been obtained, yielding single centers in at most 37 % of the irradiated spots (vs. 87 % in the present case). In the following, we demonstrate an additional step to narrow down the distribution further in case of undesired double activation, as for one of the eight sites here.
IV Controlled deactivation by photobleaching
A symmetrical approach to the activation protocol can be used for selective deactivation. B centers are very stable under laser illumination at low to moderate powers, up to about 1-2 mW, as testified by week-long measurements in PL and resonant excitation [6, 7]. This power regime is therefore considered as “safe” for the emitters. However, the B centers can photobleach when exposed to a continuous-wave (cw) PL laser at high power ( mW) during a few tens to hundreds of seconds [31]. We demonstrate how this phenomenon can be used to selectively remove an emitter from an irradiation site where two emitters are present, using cw laser drive in the high-power, “unstable” regime. We base our approach on real-time monitoring of the PL during laser illumination, in analogy with the activation case. To illustrate this technique, we consider the site of the array shown on Figure 4b (dotted orange circle) that exhibits . Figure 5a shows the of this site before selective bleaching, measured in the pulsed regime. The center peak has a value of . We apply the selective photobleaching technique while monitoring the PL signal, shown on Figure 5b. The process is as follows:
@@ IV Controlled deactivation by photobleaching
V Discussion about the conversion mechanism
@@
The ability to simultaneously resolve UV and B centers at the individual level has allowed us to observe electron-beam induced direct interconversion between these two species. Such events can shed additional light on the microscopic structure of both defects. Recent experimental studies identify the UV defect as a horizontal carbon dimer (CBCN) [21] (Fig. 6a) and the B center as a vertical carbon dimer in a split interstitial configuration (Fig. 6b), either at a boron site or at a nitrogen site [24]. In the light of these conclusions, our results can be interpreted as the observation of electron-beam induced modification of a carbon complex, and provides additional credit to both identifications. Indeed, no other pair of candidates for either UV or B centers can be easily linked together among the plausible candidates proposed in the literature [19, 32, 20, 26, 27, 21, 24]. However, a subtlety stands in the fact that the vertical dimer structure replaces only one atom of the hBN lattice, while the horizontal dimer occupies two atomic sites (Fig. 6). Therefore, the electron-beam activation process of the B center is not a simple 90∘ rotation of the in-plane carbon dimer, and necessarily involves an additional species.
@@
To account for the additional atom in the B center structure as compared with the UV center, we propose the following scenario, where a vacancy is created when switching from the horizontal to the vertical configuration of the carbon doublet.
@@
where we have supposed that the B center is at the nitrogen site, i.e. , without loss of generality since the boron site counterpart can be obtained by permuting N and B. The conversion is reversible according to our observations. However, the vacancy is known to be mobile [33, 34]. Upon diffusion (which can be induced by the electron beam [34]), the vacancy migration prevents backwards conversion and locks the defect to the state, consistently with the observed stability of the B centers at long times.
@@
Another possible scenario would involve the recombination of a boron interstitial with the UV center to form a B center: . We do not retain this scenario since it would require the unlikely presence of interstitials at a high concentration so to allow conversion of UV centers with a high yield. Additionally, since boron interstitials are mobile in hBN [33] and the UV-B conversion is reversible, the long-time stable state upon migration of the interstitial is the UV center , which is inconsistent with observations. Therefore, we believe that the first scenario describing the dissociation of the in-plane carbon dimer into a vertical dimer and a vacancy, followed by migration of the vacancy, explains both the partial reversibility at short times, and the long-time stability. Furthermore, due to the sizably higher mobility of the boron vacancy as compared with the nitrogen vacancy [35, 33], we favor the choice of rather than as the final product of the electron-beam-induced conversion.
VI Conclusion

### v2

@@ Abstract
Among the variety of quantum emitters in hexagonal boron nitride (hBN), blue-emitting color centers, or B centers, have garnered a particular interest owing to their excellent quantum optical properties. Moreover, the fact that they can be locally activated by an electron beam makes them suitable for top-down integration in photonic devices. However, in the absence of a real-time monitoring technique sensitive to individual emitters, the activation process is stochastic in the number of emitters, and its mechanism is under debate. Here, we implement an in-situ cathodoluminescence monitoring setup capable of detecting individual quantum emitters in the blue and ultraviolet (UV) range. We demonstrate that the activation of individual B centers is spatially and temporally correlated with the deactivation of individual UV centers emitting at 4.1 eV, which are ubiquitous in hBN. We then make use of the ability to detect individual B center activation events to demonstrate the controlled creation of an array with only one emitter per irradiation site. Additionally, we demonstrate a symmetric technique for heralded selective deactivation of individual emitters. Our results provide insights into the microscopic structure and activation mechanism of B centers, as well as versatile techniques for their deterministic integration.
Keywords: hexagonal boron nitride; color centers; cathodoluminescence; 2D materials; single-photon emission; quantum photonics
@@
Optically active point defects in crystals, commonly referred to as color centers, are of great significance in quantum information science due to their ability to act as single-photon emitters (SPEs) and their compatibility with nanostructures and devices.2 A well-known example is the nitrogen-vacancy (NV) center in diamond, which forms the foundation of many emerging quantum network architectures.3 More recently, the discovery of color centers in van der Waals materials has opened new avenues for integration and applications in quantum technologies,4 with hexagonal boron nitride (hBN) being a particularly promising host. This wide-bandgap semiconductor can indeed integrate high quality SPEs.5, 6 A specific class of hBN emitters is distinguished by its remarkable quantum optical properties. The so-called B centers reproducibly emit at 436 nm and exhibit photon indistinguishability at low temperature.7, 8 They are not found natively but can be locally activated by a focused electron beam,9, 10, 11 which is an asset for the controlled integration in photonic devices.12, 13 However, the activation mechanism is still debated and yields an unpredictable number of SPEs, due to the lack of a technique for heralding the successful creation of single emitters.
@@
The microscopic structure of the B centers is still under debate. Some important clues have been however identified. In particular, the B center activation is efficient only in crystals exhibiting the optical signature of UV centers.11 These UV color centers emitting at 4.1 eV are commonly found in carbon-rich hBN14, 15, 16, 17, 18, 19 and are attributed to in-plane carbon dimers (CBCN).20, 21, 22 Cathodoluminescence (CL) has been shown to allow for real-time monitoring of the emitter population dynamics under electron irradiation, allowing to study the kinetics of B center creation23 and to observe the concurrence of a decrease in the population of UV centers,24 albeit with a different timescale. These studies have been performed on ensembles and at room temperature, therefore the possible relation between UV and B centers remains unclear. Moreover, the activation and deactivation of presumably individual B centers has been observed,9, 25 but without insights about the mechanism of such dynamics and the relation to individual UV centers. Concerning the microscopic structure of B centers, several candidates have been proposed, such as a split nitrogen interstitial,26, 27 a carbon tetramer,28, 29 and a vertical carbon dimer.26 The latter hypothesis seems to have been confirmed very recently by a bundle of consistent clues obtained with combined STEM/HAADF and CL measurements.25
@@
Here, we investigate the activation and deactivation of individual B and UV centers upon irradiation by an electron beam, through a specifically designed CL detection setup based on avalanche photodiodes (APDs), which makes the in-situ monitoring sensitive to the single emitter level in both wavelength ranges. We show that individual creation events of B centers are accompanied by a simultaneous abrupt deactivation of a UV center, correlated both in time and space. This suggests a one-to-one transformation of UV to B centers. The conversion is partially reversible, as the opposite process is also observed. At long times, B centers stabilize into a steady state population, as demonstrated in prior work.23, 24 In a second step, we make use of the in-situ monitoring signal as a marker that heralds successful creation of individual emitters. We demonstrate a small array of emitters with strongly sub-Poissonian statistics of emitter numbers, as confirmed with the measurement of their second-order correlation function. We then show that accidental double creations can be corrected for using heralded photobleaching, further reducing the variations in emitter numbers. In a conclusive discussion, we tentatively provide a mechanism for supporting the observed UV-blue conversion, consistent with the vertical carbon dimer hypothesis for the microscopic structure of B centers.
Interconversion of UV and B centers
@@
We investigate hBN exfoliated from high-pressure, high-temperature grown crystals natively incorporating carbon impurities (see Methods for sample preparation). The experimental setup which allows for in-situ monitoring of the activation and deactivation of individual color centers in the UV and blue spectral range is presented in Figure 1a. The sample is placed inside a scanning electron microscope (SEM), where a holed parabolic mirror is positioned above the sample to collect the emitted photons outside the microscope. A spectrometer is placed at the output port of the SEM. Figure 1b shows the CL spectrum of a hBN crystal under electron irradiation. Three recognizable features can be observed. The exciton emission is visible at about 215 nm; the three equally spaced peaks in the 300-350 nm range are the signature of the UV emitter CBCN with a zero-phonon line at 305 nm and its phonon replica; the peaks at 440 nm and 465 nm originate respectively from the B center zero-phonon line and optical phonon replica.14, 30, 22 Using a flip-mount mirror, we can channel the emission to a single-photon detection system. The emitted light is then separated into two frequency bands, UV and blue regions respectively, using a dichroic mirror and bandpass filters. The bandwidth of both detection channels is depicted by the shaded areas on Fig. 1b. Both signals are simultaneously detected by APDs. This setup allows to be sensitive to the CL signal associated with individual UV and B centers.
@@
We first select a hBN crystal of thickness 145 nm, where the density of UV centers is low enough to spatially resolve individual emitters. Figure 2a shows a series of four consecutive CL scans of the same region of space. The map size is m m, corresponding to pixels. The scanning time is about ten minutes, and the scan direction is from left to right, then from top to bottom. Both UV and blue channels are recorded in parallel, simultaneously. The top (bottom) panels plots the UV (blue) band signal. We refer to these successive maps as scan 1 to 4. In both spectral ranges, various point emitters can be observed. In addition, some emitters of both species either appear or disappear during scans. To ease the discussion, we select five representative sites, depicted by circles in both spectral regions and labeled by letters A to E. The circle colors indicate that the emitters are active (green circles), inactive (red circles) or switching (two-colored circles). In the UV maps, sites A and B indicate UV emitters that are active during scans 1 and 2 (upper panel), while no blue emitter is observed at these same sites (lower panel). In scan 3, the UV emitters located sites A and B are deactivated, which shows up as interrupted spots with CL signal only in the upper part. Simultaneously, signal appears in the blue range at the same locations. The switching appears instantaneous at the timescale of the integration time per pixel (100 ms), as the blue signal appears at the same pixels where the UV signal disappears. In scan 4, no UV emitter is visible in sites A and B, while blue emitters now appear as full spots. Sites C and D exhibit the inverse process along the four scans, with blue emitters initially present in scan 1, switching to UV during scan 2, and staying visible as UV emitters in the two following scans. Finally, site E undergoes a bidirectional process, with an initial switching from UV to blue during scan 1, staying in the blue state during scans 2 and 3, and switching back to UV during scan 4. Other similar events can be observed on these CL scans, as well as all other scans we performed. We illustrate this further by showing consecutive scans of two emitters on Figure 2b and c. Both emitters exhibit multiple switching events between the UV and the blue state while the electron beam is passing over them. Additional examples of consecutive scans are shown in the Supporting Information section S1. We note that the irradiation parameters are chosen so that only a few emitters switch at every scan, so to make the observation of the interconversion more evident. Since it is a stochastic process, some UV emitters are expected to persist during many frames. However, it cannot be excluded that some of the persistent spots observed on the UV maps originate from other defect species. CL spectra suggest however that the UV emission is largely dominated by contribution from the carbon dimers, and therefore that most of the UV signal originates from UV centers. On the other hand, in the blue wavelength range, regions that have not been exposed to the electron beam are initially devoid of CL signal, so that all spots are attributed to B centers (see Supporting Information section S1). Overall, our observations can be summarized as follows.
@@
(4) The switching occurs exactly under the electron beam, as testified by truncated CL spots, with the signal appearing or disappearing at the emitter position. This can be observed as incomplete lines of bright pixels switching while the electron beam is scanning over the SPE. This implies that the process is driven locally by the electron beam, at the immediate vicinity of the UV centers. This also demonstrates that, in samples with high concentration of UV centers, the resolution of the B center creation is solely given by the SEM resolution.
(5) Essentially all occurrences where a single UV center disappears while scanning are associated with the appearance of a B center. This suggests that the conversion to B center is the main – if not the only – decay process of the UV center under electron irradiation. Other mechanisms were proposed, which consist in the dissociation of the CBCN dimers into (CB,CN) donor-acceptor pairs at various neighboring positions.31 Given that such mechanisms would lead to the disappearance of UV centers without simultaneous activation of a B center, which we do not observe in our datasets, we conclude that they are likely slower or less probable, at least in our range of irradiation parameters.
@@
Heralded activation of B centers
We now select a flake with a higher density of UV centers, such that conversion can occur at any location. Such flakes exhibit a large UV signal, similar as the spectrum shown Fig. 1b. We continuously irradiate the crystal at a fixed position, while recording the signal of B centers with the APD. Figure 3 shows a representative example of the intensity trace measured during 80 s. Discrete increases of the signal can be observed, showing that individual creation events can be unambiguously identified. The signal associated with the creation of individual emitters is about to counts per second. A few downward steps are also visible, associated with deactivations as observed in the previous section. However, the overall trend is a global increase of the number of emitters, consistently with prior work on large ensembles.23, 24 This suggests that the blue emitters are stabilized under continuous irradiation. Each upward step constitutes a signal heralding the success of individual B center activation.
@@
The sample is subsequently characterized in photoluminescence (PL) at room temperature, in a confocal microscope (see Methods). A pulsed laser emitting at 405 nm is used to excite the emitters. Figure 4b shows a PL confocal map of the sample. Eight luminescent spots are visible. The slight misalignment of the spot array originates from hysteresis of the piezoelectric positioner used to define the irradiation sites. The emitters are characterized in a Hanbury Brown and Twiss setup to estimate the number of emitters. The second-order correlation function is measured as the histogram of the delay time between consecutive photon detection events, and fitted using a multipeak function (see Methods). A value of is attributed to a single emitter. An example of is shown on Figure 4c, corresponding to the circled spot on Figure 4b. Among the eight sites, seven are single emitters, and one, indicated by the dotted circle on Figure 4b, has a higher . The latter corresponds to an event where two steps were observed before the electron beam was blanked. A workaround for such cases is proposed in the next section.
This proof of principle demonstrates the ability to control the number of activated color centers. Without the active feedback developed in this work using in-situ CL monitoring, a Poisson distribution of the emitter numbers would have been obtained, yielding single centers in at most 37 % of the irradiated spots, which is the theoretical higher bound allowed by the Poisson distribution. We note that, in the absence of in-situ monitoring, a yield of 33 % has been obtained using calibrated irradiation.11 In the following, we demonstrate an additional step to narrow down the distribution further in case of undesired double activation, as for one of the eight sites here.
Controlled deactivation by photobleaching
A symmetrical approach to the activation protocol can be used for selective deactivation. B centers are very stable under laser illumination at low to moderate powers, up to about 1-2 mW, as testified by week-long measurements in PL and resonant excitation.7, 8 This power regime is therefore considered as “safe” for the emitters. However, the B centers can photobleach when exposed to a continuous-wave (cw) PL laser at high power ( mW) during a few tens to hundreds of seconds.32 We demonstrate how this phenomenon can be used to selectively remove an emitter from an irradiation site where two emitters are present, using cw laser drive in the high-power, “unstable” regime. We base our approach on real-time monitoring of the PL during laser illumination, in analogy with the activation case. To illustrate this technique, we consider the site of the array shown on Figure 4b (dotted orange circle) that exhibits . Figure 5a shows the of this site before selective bleaching, measured in the pulsed regime. The center peak has a value of . We apply the selective photobleaching technique while monitoring the PL signal, shown on Figure 5b. The process is as follows:
@@
Discussion about the conversion mechanism
@@
The ability to simultaneously resolve UV and B centers at the individual level has allowed us to observe electron-beam induced direct interconversion between these two species. Such events can shed additional light on the microscopic structure of both defects. Recent experimental studies identify the UV defect as a horizontal carbon dimer (CBCN)22 (Fig. 6a) and the B center as a vertical carbon dimer in a split interstitial configuration (Fig. 6b), either at a boron site or at a nitrogen site .25 These proposed atomic structures for both species have also been investigated using density functional theory (DFT) calculations,20, 21, 25 giving additional credit to these structural models. In the light of these conclusions, our results can be interpreted as the observation of electron-beam induced modification of a carbon complex, and provides additional credit to both identifications. Indeed, no other pair of candidates for either UV or B centers can be easily linked together among the plausible candidates proposed in the literature.20, 33, 21, 27, 28, 22, 25 However, a subtlety stands in the fact that the vertical dimer structure replaces only one atom of the hBN lattice, while the horizontal dimer occupies two atomic sites (Fig. 6). Therefore, the electron-beam activation process of the B center is not a simple 90∘ rotation of the in-plane carbon dimer, and necessarily involves an additional species.
@@
To account for the additional atom in the B center structure as compared with the UV center, we tentatively propose the following scenario, where a vacancy is created when switching from the horizontal to the vertical configuration of the carbon doublet.
@@
where we have supposed that the B center is at the nitrogen site, i.e. , without loss of generality since the boron site counterpart can be obtained by permuting N and B. The conversion is reversible according to our observations. However, the vacancy is known to be mobile.34, 35 Upon diffusion (which can be induced by the electron beam35), the vacancy migration prevents backwards conversion and locks the defect to the state, consistently with the observed stability of the B centers at long times.
@@
Another possible scenario would involve the recombination of a boron interstitial with the UV center to form a B center: . We do not retain this scenario since it would require the unlikely presence of interstitials at a high concentration so to allow conversion of UV centers with a high yield. Additionally, since boron interstitials are mobile in hBN34 and the UV-B conversion is reversible, the long-time stable state upon migration would be the UV center , which is inconsistent with observations. Therefore, we believe that the first scenario describing the dissociation of the in-plane carbon dimer into a vertical dimer and a vacancy, followed by migration of the vacancy, explains both the partial reversibility at short times, and the long-time stability. Furthermore, due to the sizably higher mobility of the boron vacancy as compared with the nitrogen vacancy,36, 34 we favor the choice of rather than as the final product of the electron-beam-induced conversion.
Conclusion
We have observed electron-beam induced interconversion between individual UV centers and B centers, appearing as abrupt anticorrelated activation or deactivation events of the point defect emission on simultaneously recorded CL scans. The sensitivity at the single-emitter level, combined with the spatial resolution required to track individual emitters, makes it possible to overcome the limitations inherent to ensemble CL measurements. In ensemble experiments such as those realized in prior work,23, 24, 31 the CL signal is generally not directly proportional to the actual defect concentration due to irradiation-induced modifications of the excitation and collection efficiencies during the measurement. Such effects can hinder the direct observation of one-to-one interconversion between B and UV centers, as we discuss in the Supporting Information section S3.